# Gautschi-type and implicit–explicit integrators for constrained wave equations

### v1

@@ Abstract.
This paper deals with the construction and analysis of two integrators for (semi-linear) second-order partial differential–algebraic equations of semi-explicit type. More precisely, we consider an implicit–explicit Crank–Nicolson scheme as well as an exponential integrator of Gautschi type. For this, well-known wave integrators for unconstrained systems are combined with techniques known from the field of differential–algebraic equations. The result are efficient time stepping schemes, which are provable of second order. Moreover, we discuss the practical implementation of the Gautschi-type method, which involves the solution of certain saddle point problems. The theoretical results are verified by numerical experiments for the the wave equation with kinetic boundary conditions.
@@ 1. Introduction
with initial conditions $u(0)=u^{0}$, $\dot{u}(0)=w^{0}$ which underlie an additional (linear) constraint of the form $\mathcal{B}u(t)=g(t)$. Such systems can be considered as differential–algebraic equations (DAEs) in Banach spaces, also called partial differential–algebraic equations (PDAEs); see [LMT13, Alt15] for an introduction. Second-order PDAEs of wave type appear, e.g., in the field of elastic multibody dynamics [Sim98, Sim13], in problems on wave propagation with non-standard boundary conditions [Alt23], or, more generally, if two second-order systems are coupled.
@@
For the time discretization of (1.1) (without a constraint) the two main classes of methods are given by explicit and implicit schemes. The advantage of an explicit time-stepping lies in the easy implementation and possibility for parallelization. However, when combined with a spatial discretization small time steps have to be chosen. On the other hand, implicit methods are unconditionally stable and thus allow for large time steps. Here, however, a nonlinear system of equations has to be solved in each time step. Thus, it is desirable to use a scheme that takes the best from both methods and obtain a scheme which is unconditionally stable in the linear part, and treats the nonlinear term explicitly. In this work, we consider two types of methods, namely an implicit–explicit Crank–Nicolson scheme (IMEX) as well as an exponential integrator of Gautschi type.
@@
The idea of IMEX schemes is to split the equation in a linear unbounded part that is treated implicitly and a (possibly nonlinear) part which is treated explicitly, under the condition that the scheme is stable under a step-size restriction which only depends on the Lipschitz condition of the explicit part and preserves the order of convergence of the underlying schemes. For IMEX schemes, there is a well-developed theory which covers IMEX multistep schemes [ARW95, FHV97, ACM99, HR07] as well as IMEX Runge–Kutta methods [ARS97, Bos07]. More recently, an IMEX scheme was proposed which exploits the structure of the second-order system by a combination of the leapfrog method with the Crank–Nicolson method [HL21]. This scheme will be the basis of our IMEX method for the constrained problem. Further, there is the so-called Crank–Nicolson–leapfrog scheme proposed in [LT12, LLT16], which uses the leapfrog method for the first-order system, and is stable only when the explicitly treated part is skew-symmetric.
@@
In the undamped case $\mathcal{D}=0$, also Gautschi-type integrators provide a powerful tool for the time integration of wave-type equations. These methods follow the methodology of exponential integrators in the parabolic setting [HO10]. Hence, the idea is to integrate the linear (stiff) part in an exact manner, which allows large time steps that are not restricted by the frequencies of the differential operator. For second-order wave-type systems (without a constraint), Gautschi-type integrators were developed in [Gau61, HL99, Gri02]. For this, one considers the first-order formulation and applies the variation-of-constants formula. The combination of the resulting formula for times $t+\tau$ and $t-\tau$ and an approximation of the integral term by the simplest left/right rectangle rule finally yields an efficient two-step time integration scheme. In the same way as exponential integrators rely on an efficient computation of the matrix exponential applied to a vector [MV78], Gautschi-type integrators are based on the cosine of a matrix (applied to a vector). Also this can be realized efficiently [HAM10], e.g., by using Krylov subspace methods [EE06, CKN24].
@@
In this paper, we aim to translate these two concepts to wave-type equations (1.1) under a linear constraint. For the construction of our novel PDAE integrators, we decompose the solution as in [EM13]. This means that we consider the part of the solution, which lies in the kernel of the constraint operator, separately. For the IMEX scheme, we use the half-step formulation from [HL21] and impose the linear constraints only on the time derivative of the solution such that the solution $u$ satisfies the constraint exactly. This leads to a method which requires (after the spatial discretization) the solution of two (linear) saddle point problems in each time step, one including the mass and one the stiffness matrix.
@@
For the Gautschi-type integrators, we again consider only the part of the solution in the kernel, which together with techniques developed for constrained parabolic systems in [AZ20], then leads to the desired explicit integration scheme. To prove (mesh-independent) convergence of this novel integrator, the theory of [HL99, Gri02] for the unconstrained case is combined with the already mentioned PDAE techniques. In view of the efficiency of the practical implementation, the critical part is the computation of $\cos(\tau\Omega_{\ker})v$ for some vector $v$. Here, $\Omega_{\ker}$ denotes the root of the differential operator $\mathcal{A}$ (or its discrete counterpart) restricted to the kernel of the constraint operator $\mathcal{B}$. For this, we propose an Arnoldi method, which does not require the computation of $\Omega_{\ker}$; cf. [Gri05] for the unconstrained case. Moreover, by particularly constructed saddle point problems, the computations can be restricted to computations on the kernel of the constraint operator. At this point, we would like to emphasize that Gautschi-type integrators are tailored to problems with vanishing damping, i.e., $\mathcal{D}=0$. In order to include damping in an exponential manner, one could also use general exponential methods of Runge–Kutta or multistep type; cf. the review [HO10]. This, however, is beyond the scope of this work.
@@
The paper is organized as follows. In Section 2 we introduce the system equations in detail and discuss its unique solvability. Afterwards, we turn to the numerical treatment of theses systems. First, an implicit–explicit version of the Crank–Nicolson scheme is introduced and analyzed in Section 3. Here, a detailed inspection of all the error contributions allows us to prove an error equation which can be resolved by the techniques developed for the unconstrained IMEX scheme. It follows the discussion of Gautschi-type integrators for damping-free systems in Section 4. Besides proving second-order accuracy of the method, we also provide comments on the practical implementation of the algorithm. Finally, we present a numerical example in Section 5 and conclude in Section 6.
@@ 2. Wave-type Systems with Linear Constraints
This section is devoted to the introduction of semi-linear PDAEs of wave type. Hence, we consider systems of the form (1.1) which underlie an additional (linear) constraint.
@@ 2.1. Problem formulation
In this subsection, we introduce a class of hyperbolic wave-type systems, which are restricted by a linear constraint. Allowing (sufficiently smooth) nonlinearities in the low-order terms of the dynamic equation, we end up with the following PDAE: find abstract functions $u\colon[0,T]\to\mathcal{V}$ and $\lambda\colon[0,T]\to\mathcal{Q}$ such that
@@
(2.1b) $\mathcal{B}u(t)$ $=g(t)$ in $\qquad\text{in }\mathcal{Q}^{*}.$
Therein, the spaces $\mathcal{V}$ and $\mathcal{Q}$ are assumed to be Hilbert spaces with respective duals $\mathcal{V}^{*}$ and $\mathcal{Q}^{*}$. Further, the space $\mathcal{V}$ is part of a Gelfand triple $\mathcal{V}$, $\mathcal{H}$, $\mathcal{V}^{*}$, i.e., we assume a continuous (and dense) embedding $\mathcal{V}\hookrightarrow\mathcal{H}$, which in turn implies a corresponding embedding of the duals [Zei90, Ch. 23.4].
For second-order systems in the weak form, one typically seeks a solution with the first and second derivative taking values in $\mathcal{H}$ and $\mathcal{V}^{*}$, respectively. Hence, initial data is expected to be of the form $u(0)=u^{0}\in\mathcal{V}$, $\dot{u}(0)=w^{0}\in\mathcal{H}$. Note, however, that the constraint (2.1b) typically implies certain consistency conditions on the initial data; cf. [Alt15, Rem. 7.4]. At this point, it is reasonable to assume $u^{0},w^{0}\in\mathcal{V}$ with $\mathcal{B}u^{0}=g(0)$ and $\mathcal{B}w^{0}=\dot{g}(0)$.
For the linear operators $\mathcal{A}\in\mathcal{L}(\mathcal{V},\mathcal{V}^{*})$, $\mathcal{B}\in\mathcal{L}(\mathcal{V},\mathcal{Q}^{*})$, and $\mathcal{D}\in\mathcal{L}(\mathcal{H},\mathcal{V}^{*})$ appearing in (2.1), we make the following assumptions.
@@
The operator $\mathcal{B}\colon\mathcal{V}\to\mathcal{Q}^{*}$ is linear, continuous, and satisfies an inf-sup condition, i.e., there exists a constant $\beta>0$ such that
@@
The operator $\mathcal{A}\colon\mathcal{V}\to\mathcal{V}^{*}$ is linear, continuous, and elliptic on $\mathcal{V}_{\ker}\coloneqq\ker\mathcal{B}$, i.e., on the kernel of the constraint operator. Furthermore, we assume that we can decompose the operator into $\mathcal{A}=\mathcal{A}_{1}+\mathcal{A}_{2}$, with $\mathcal{A}_{1}$ being self-adjoint and $\mathcal{A}_{2}\in\mathcal{L}(\mathcal{V},\mathcal{H}^{*})$. In particular, we may assume that $\mathcal{A}_{1}$ is elliptic on $\mathcal{V}_{\ker}$ as well; cf. [AZ20].
@@
(2.2) $(\mathcal{D}u,u)_{\mathcal{H}}\geq 0\quad\text{ for all }\quad u\in\mathcal{V}.$
@@
The previous assumptions can be generalized to the case where $\mathcal{A}$ only satisfies a Gårding inequality, see [AZ20], or $\mathcal{D}$ satisfies a lower bound of the type $(\mathcal{D}u,u)_{\mathcal{H}}\geq-c_{\mathcal{D}}\|u\|^{2}_{\mathcal{H}}$.
As discussed in [AZ20], Assumption 2.1 implies the existence of a continuous right-inverse $\mathcal{B}^{-}\colon\mathcal{Q}^{*}\to\mathcal{V}$, which leads to the decomposition
(2.3) $\mathcal{V}=\mathcal{V}_{\ker}\oplus\mathcal{V}_{\text{c}}\qquad\text{with}\qquad\mathcal{V}_{\ker}=\ker\mathcal{B},\quad\mathcal{V}_{\text{c}}=\operatorname{im}\mathcal{B}^{-}.$
Following [AZ18], we define the complementary space $\mathcal{V}_{\text{c}}$ as
@@
Here, we use the fact that functionals in $\mathcal{V}^{*}$ define functionals in $\mathcal{V}_{\ker}^{*}$ simply through the restriction to $\mathcal{V}_{\ker}$. Note that $\mathcal{A}_{\ker}$ is elliptic due to Assumption 2.2. Hence, $-\mathcal{A}_{\ker}$ generates an analytic semigroup on the closure of $\mathcal{V}_{\ker}$ in the $\mathcal{H}$-norm, i.e., on $\mathcal{H}_{\ker}\coloneqq\operatorname{clos}_{\mathcal{H}}\mathcal{V}_{\ker}$; cf. [Paz83, Ch. 7, Th. 2.7].
@@
With the space $\mathcal{H}_{\ker}$ at hand, the assumptions on the initial data may be relaxed to $w^{0}\in\mathcal{H}_{\ker}+\mathcal{B}^{-}\dot{g}(0)$, which is still consistent with the constraint.
@@
One of the following two conditions is satisfied:
@@
Consider the wave equation with additional convection and damping terms,

### v2

@@ Abstract.
This paper deals with the construction and analysis of two integrators for (semi-linear) second-order partial differential–algebraic equations of semi-explicit type. More precisely, we consider an implicit–explicit Crank–Nicolson scheme as well as an exponential integrator of Gautschi type. For this, well-known wave integrators for unconstrained systems are combined with techniques known from the field of differential–algebraic equations. This results in efficient time stepping schemes that are provable of second order. Moreover, we discuss the practical implementation of the Gautschi-type method, which involves the solution of certain saddle point problems. The theoretical results are verified by a numerical experiment for the wave equation with kinetic boundary conditions.
@@ 1. Introduction
with initial conditions $u(0)=u^{0}$, $\dot{u}(0)=w^{0}$, which underlie an additional (linear) constraint of the form $\mathcal{B}u(t)=g(t)$. Typically, $\mathcal{A}$ and $\mathcal{D}$ are differential operators; see Examples 2.8 and 2.9 below. Such systems can be considered as differential–algebraic equations (DAEs) in Banach spaces, also called partial differential–algebraic equations (PDAEs); see [LamMT13, Alt15] for an introduction. Second-order PDAEs of wave type appear, e.g., in the field of elastic multibody dynamics [Sim98, Sim13], in problems on wave propagation with non-standard boundary conditions [Alt23], or, more generally, if two second-order systems are coupled.
@@
For the time discretization of (1.1) (without a constraint) the two main classes of methods are given by explicit and implicit schemes. The advantage of an explicit time-stepping lies in the easy implementation and possibility for parallelization. However, when combined with a spatial discretization small time steps have to be chosen. On the other hand, implicit methods are unconditionally stable and thus allow for large time steps. Here, however, a nonlinear system of equations has to be solved in each time step. Thus, it is desirable to use a scheme that takes the best from both methods and obtain a scheme that is unconditionally stable in the linear part, and treats the nonlinear term explicitly. In this work, we consider two types of methods, namely an implicit–explicit Crank–Nicolson scheme (IMEX) as well as an exponential integrator of Gautschi type.
@@
The idea of IMEX schemes is to split the equation in a linear unbounded part that is treated implicitly and a (possibly nonlinear) part treated explicitly, under the condition that the scheme is stable under a step-size restriction that only depends on the Lipschitz condition of the explicit part and preserves the order of convergence of the underlying schemes. For IMEX schemes, there is a well-developed theory, covering IMEX multistep schemes [AscRW95, FraHV96, AkrCM99, HunR07] as well as IMEX Runge–Kutta methods [AscRS97, Bos07]. More recently, an IMEX scheme was proposed that exploits the structure of the second-order system by a combination of the leapfrog method with the Crank–Nicolson method [HocL21]. This scheme will be the basis of our IMEX method for the constrained problem. Further, there is the so-called Crank–Nicolson–leapfrog scheme proposed in [LayT12, LayLR16], which uses the leapfrog method for the first-order system, and is stable only when the explicitly treated part is skew-symmetric.
@@
In the undamped case $\mathcal{D}=0$, also Gautschi-type integrators provide a powerful tool for the time integration of wave-type equations. These methods follow the methodology of exponential integrators in the parabolic setting [HocO10]. Hence, the idea is to integrate the linear (stiff) part in an exact manner, which allows large time steps that are not restricted by the frequencies of the differential operator. For second-order wave-type systems (without a constraint), Gautschi-type integrators were developed in [Gau61, HocL99, Gri02]. For this, one considers the first-order formulation and applies the variation-of-constants formula. The combination of the resulting formula for times $t+\tau$ and $t-\tau$ and an approximation of the integral term by the simplest left/right rectangle rule finally yields an efficient two-step time integration scheme. In the same way as exponential integrators rely on an efficient computation of the matrix exponential applied to a vector [MolV78], Gautschi-type integrators are based on the cosine of a matrix (applied to a vector). Also this can be realized efficiently [HigA10], e.g., by using Krylov subspace methods [EieE06, CorKN24].
@@
For general constrained systems, one often considers algebraically stable Runge–Kutta methods [HaiLR89, KunM06, AltZ18]. These methods, however, do not take advantage of the given saddle point structure. Related approaches for systems of first order include splitting methods [AltO17], (discontinuous) Galerkin methods [VouR18, AltH21], as well as exponential integrators [AltZ20]. In this paper, we aim to translate the previously mentioned concepts to wave-type equations of the form (1.1) under a linear constraint. For the construction of our novel PDAE integrators, we decompose the solution as in [EmmM13]. This means that we consider the part of the solution that lies in the kernel of the constraint operator, separately. For the IMEX scheme, we use the half-step formulation from [HocL21] and impose the linear constraints only on the time derivative of the solution such that the solution $u$ satisfies the constraint exactly. This leads to a method that requires (after the spatial discretization) the solution of two (linear) saddle point problems in each time step, one including the mass and one the stiffness matrix.
@@
For the Gautschi-type integrators, we again consider only the part of the solution in the kernel. Together with techniques developed for constrained parabolic systems in [AltZ20], this then leads to the desired explicit integration scheme. To prove (mesh-independent) convergence of this novel integrator, the theory of [HocL99, Gri02] for the unconstrained case is combined with the already mentioned PDAE techniques. In view of the efficiency of the practical implementation, the critical part is the computation of $\cos(\tau\Omega_{\ker})v$ for some vector $v$. Here, $\Omega_{\ker}$ denotes the root of the differential operator $\mathcal{A}$ (or its discrete counterpart) restricted to the kernel of the constraint operator $\mathcal{B}$. For this, we propose an Arnoldi method that does not require the computation of $\Omega_{\ker}$; cf. [Gri05b] for the unconstrained case. Moreover, by particularly constructed saddle point problems, the computations can be restricted to computations on the kernel of the constraint operator. At this point, we would like to emphasize that Gautschi-type integrators are tailored to problems with vanishing damping, i.e., $\mathcal{D}=0$. In order to include damping in an exponential manner, one could also use general exponential methods of Runge–Kutta or multistep type; cf. the review [HocO10]. This, however, is beyond the scope of this work.
@@
The paper is organized as follows. In Section 2 we introduce the system equations in detail and discuss its unique solvability. Afterwards, we turn to the numerical treatment of theses systems. First, an implicit–explicit version of the Crank–Nicolson scheme is introduced and analyzed in Section 3. Here, a detailed inspection of all the error contributions allows us to prove an error equation, which can be resolved by the techniques developed for the unconstrained IMEX scheme. It follows the discussion of Gautschi-type integrators for damping-free systems in Section 4. Besides proving second-order accuracy of the method, we also provide comments on the practical implementation of the algorithm. Finally, we present a numerical example in Section 5 and conclude in Section 6.
@@ 2. Wave-type Systems with Linear Constraints
This section is devoted to the introduction of semi-linear PDAEs of wave type. Hence, we consider systems of the form (1.1) that underlie an additional (linear) constraint.
@@ 2.1. Problem formulation
In this subsection, we introduce a class of hyperbolic wave-type systems, restricted by a linear constraint. Allowing (sufficiently smooth) nonlinearities in the low-order terms of the dynamic equation, we end up with the following PDAE: find abstract functions $u\colon[0,T]\to\mathcal{V}$ and $\lambda\colon[0,T]\to\mathcal{Q}$ such that
@@
(2.1b) $\mathcal{B}u(t)$ $=g(t)$ $\qquad\text{in }\mathcal{Q}^{*}.$
Therein, the spaces $\mathcal{V}$ and $\mathcal{Q}$ are assumed to be Hilbert spaces with respective duals $\mathcal{V}^{*}$ and $\mathcal{Q}^{*}$. Further, the space $\mathcal{V}$ is part of a Gelfand triple $\mathcal{V}$, $\mathcal{H}$, $\mathcal{V}^{*}$, i.e., we assume a continuous (and dense) embedding $\mathcal{V}\hookrightarrow\mathcal{H}$,. This, in turn, implies a corresponding embedding of the duals [Zei90a, Ch. 23.4]. The appearing operators are linear and satisfy $\mathcal{A}\in\mathcal{L}(\mathcal{V},\mathcal{V}^{*})$, $\mathcal{B}\in\mathcal{L}(\mathcal{V},\mathcal{Q}^{*})$, and $\mathcal{D}\in\mathcal{L}(\mathcal{H},\mathcal{V}^{*})$. Moreover, equation (2.1a) includes the Lagrange multiplier $\lambda$, which is used to enforce the constraint (2.1b) following the Lagrangian approach.
For second-order systems in the weak form, one typically seeks a solution with the first and second derivative taking values in $\mathcal{H}$ and $\mathcal{V}^{*}$, respectively. Hence, initial data is expected to be of the form $u(0)=u^{0}\in\mathcal{V}$, $\dot{u}(0)=w^{0}\in\mathcal{H}$. Note, however, that the constraint (2.1b) typically implies certain consistency conditions on the initial data; cf. [Alt15, Rem. 7.4]. At this point, it is reasonable to assume $u^{0},w^{0}\in\mathcal{V}$ with $\mathcal{B}u^{0}=g(0)$ and $\mathcal{B}w^{0}=\dot{g}(0)$. For the linear operators we make the following assumptions.
@@
The operator $\mathcal{B}\colon\mathcal{V}\to\mathcal{Q}^{*}$ is linear, continuous, and satisfies an inf–sup condition, i.e., there exists a constant $\beta>0$ such that
@@
The operator $\mathcal{A}\colon\mathcal{V}\to\mathcal{V}^{*}$ is linear, continuous, and elliptic on $\mathcal{V}_{\ker}\coloneqq\ker\mathcal{B}$, i.e., on the kernel of the constraint operator. Furthermore, we assume that we can decompose the operator into $\mathcal{A}=\mathcal{A}_{1}+\mathcal{A}_{2}$, with $\mathcal{A}_{1}$ being self-adjoint, i.e., $\langle\mathcal{A}_{1}u,v\rangle=\langle u,\mathcal{A}_{1}v\rangle$, and $\mathcal{A}_{2}\in\mathcal{L}(\mathcal{V},\mathcal{H}^{*})$. In particular, we may assume that $\mathcal{A}_{1}$ is elliptic on $\mathcal{V}_{\ker}$ as well; cf. [AltZ20].
@@
[…]
@@
The previous assumptions can be generalized to the case where $\mathcal{A}$ only satisfies a Gårding inequality; see [AltZ20], or $\mathcal{D}$ satisfies a lower bound of the type $(\mathcal{D}u,u)_{\mathcal{H}}\geq-c_{\mathcal{D}}\|u\|^{2}_{\mathcal{H}}$.
As discussed in [AltZ20], Assumption 2.1 implies the existence of a continuous right-inverse $\mathcal{B}^{-}\colon\mathcal{Q}^{*}\to\mathcal{V}$. This then leads to the decomposition
(2.3) $\mathcal{V}=\mathcal{V}_{\ker}\oplus\mathcal{V}_{\text{c}}\qquad\text{with}\qquad\mathcal{V}_{\ker}=\ker\mathcal{B},\quad\mathcal{V}_{\text{c}}=\operatorname{im}\mathcal{B}^{-},$
where we denote by $\oplus$ the direct sum of two vector spaces. This means, in particular, that $\mathcal{V}_{\ker}\cap\mathcal{V}_{\text{c}}=\{0\}$. Following [AltZ18], we define the complementary space $\mathcal{V}_{\text{c}}$ as
@@
Here, we use the fact that functionals in $\mathcal{V}^{*}$ define functionals in $\mathcal{V}_{\ker}^{*}$ simply through the restriction to $\mathcal{V}_{\ker}$. Note that $\mathcal{A}_{\ker}$ is elliptic due to Assumption 2.2. Hence, $\mathcal{A}_{\ker}$ generates an analytic semigroup on the closure of $\mathcal{V}_{\ker}$ in the $\mathcal{H}$-norm, i.e., on $\mathcal{H}_{\ker}\coloneqq\operatorname{clos}_{\mathcal{H}}\mathcal{V}_{\ker}$; cf. [Paz83, Ch. 7, Th. 2.7].
@@
With the space $\mathcal{H}_{\ker}$ at hand, the assumptions on the initial data may be relaxed to $w^{0}\in\mathcal{H}_{\ker}+\mathcal{B}^{-}\dot{g}(0)$, i.e., $w^{0}$ can be written as the sum of $\mathcal{B}^{-}\dot{g}(0)\in\mathcal{V}_{\text{c}}$ and an element of $\mathcal{H}_{\ker}$, which is still consistent with the constraint.
@@
Let $f$ satisfy the classical Carathéodory condition and one of the following two conditions:
@@
In this first example, we focus on the assumptions in the non-constrained part of the problem. Thus, consider the wave equation with additional convection and damping terms,
@@
Example 2.9.
In our second example, we focus on the constraint. For this, we consider
$\ddot{u}-\Delta u=f(t)\qquad\text{in }\Omega\subseteq\mathbb{R}^{3}$
subject to $\mathcal{B}u\coloneqq\operatorname{trace}u=g\in H^{1/2}(\Gamma)$ for $\Gamma\coloneqq\partial\Omega$. Here, we have $\mathcal{V}=H^{1}(\Omega)$ and $\mathcal{H}=L^{2}(\Omega)$ as well as $\mathcal{Q}=H^{-1/2}(\Gamma)$. Thus, we obtain $\mathcal{V}_{\ker}=H^{1}_{0}(\Omega)$ and $\mathcal{H}_{\ker}=\mathcal{H}$. As a result, the operator $\mathcal{A}_{\ker}$ equals the Dirichlet Laplacian and is hence elliptic on $\mathcal{V}_{\ker}$.
A more sophisticated example of a wave system with kinetic boundary conditions is postponed to Section 5, where we present a numerical example.